# Grounded Misunderstandings in Asymmetric Dialogue: A Perspectivist Annotation Scheme for MapTask

### v1

@@ Abstract
| Nan Li, Albert Gatt, Massimo Poesio |
| Utrecht University, Utrecht, The Netherlands |
@@ 1. Introduction
RQ1 Is it possible to develop a personal-interpretation annotation scheme capture nuanced understanding states for REs in collaborative tasks like MapTask?
@@
We operationalize reference expression grounding as a graded, perspective-dependent alignment process, and propose a annotation scheme at the RE level.
@@ 2. Related work
Formal dialogue theories have modeled interlocutors’ intentions and repairs through structured information state models. Some proposals Poesio and Traum (1997); Matheson et al. (2000) extended Discourse Representation Theory to integrate semantic and pragmatic information, showing how turn-taking, discourse segmentation, and grounding are represented through conversation acts in common ground. Ginzburg (2012) proposed the KoS framework111A term loosely connected to conversation-oriented semantics Ginzburg (2012)., where unresolved Questions Under Discussion (QUDs) persist until explicitly addressed. Lascarides and Asher (2009) modeled agreement and correction dynamics but assumed a highly idealized scenario which excludes misunderstandings. Schlangen and Skantze (2011) treated divergent understandings as a sub-utterance level hypothesis-revision process. Some recent work leveraged annotations from other meaning represenatations e.g., Abstract Meaning Representation, to represent common ground structure and information states Khebour et al. (2024); Lai et al. (2025). We are informed by this line of work, but focus on RE-level perspectivist interpretations and inferring understanding states through the alignment and misalignment of these interpretations.
@@ 3.1. Development and Follow-up Studies
HCRC MapTask is a corpus documenting unscripted dialogues in a collaborative route-following task with asymmetric maps. Participants must collaborate to reproduce a route on one participant’s map (the instruction giver’s) on the other’s (the instruction follower’s). They cannot see each other’s maps and can only communicate verbally, with eye contact possible in half of the dialogues. The intentional differences between maps on routes, landmark names, and landmark placements often leads to ambiguous or misleading instructions. This makes the MapTask corpus a valuable resource and testbed for studying grounding, reference, and repair theories in dialogue.
Subsequent research analyzed reference expressions in MapTask from multiple perspectives. Bard et al. (2000) show that speakers reduce reduce articulatory clarity when mentioning landmarks for the second time, even when the addressee has not heard the first mention or cannot see the referent, suggesting that intelligibility control is driven primarily by the speaker’s own knowledge rather than careful modeling of the listener’s information state. Varges (2005) treated spatial descriptions in MapTask as reference expressions that distinguish particular map points from distractors, observing that speakers frequently use spatial relations (e.g., “above the west lake”) and vague quantifiers (e.g., “about two inches southwest”) when referring to featureless route points around named landmarks. Marge and Rudnicky (2011) analyzed miscommunication patterns in MapTask, documenting how instruction followers detect ambiguity (e.g., multiple possible referents) and misunderstanding through clarification questions, repair strategies, and explicit feedback when requested actions appear impossible or unclear. Poesio et al. (2004) applied the MATE annotation scheme Poesio et al. (1999) to an Italian adaptation of MapTask to make anaphoric relations explicit.
@@ 4.1. Approach and General Principles
ordinal: Positional index (0, 1, …) for landmarks of the same type occurring multiple times on maps, ordered from bottom to top. Landmarks of this type have the same ordinal number if they are at the same location on both maps, which means it is possible that the only “parked van” on the follower’s map is m0_parked_van#1@f if the giver’s map has two “parked van”, while on the giver’s map m0_parked_van#0@g denotes the unique lower instance.
@@ 4.2. Annotation Pipeline
The prompt (see Figure 2) guides the model through a sequential and heuristic annotation workflow. First, the model identifies the speaker and addressee roles (giver/follower) and assigns the speaker’s intended landmark ID. Second, it evaluates is_quantificational to determine whether the expression presupposes a referent. If not quantificational, the model proceeds through the remaining four attributes in order. Finally, if is_grounded=true, the model assigns the addressee’s interpreted landmark ID. The model is asked to give a concise textual reason citing dialogue evidence for our further analysis.
@@ 4.3. Quality Evaluation
We apply the annotation pipeline to all 128 dialogues in the MapTask corpus, resulting in 13,077 annotated reference expression instances 5554 reference expressions (from 3 dialogues) were not annotated by the LLM despite being present in the dialogue context and RE list, likely due to uncertainty in generation. We will verify these cases in follow-up runs..
To evaluate the reliability of the annotation pipeline, we selected 3 dialogues (dialogue IDs: q1ec2, q1nc3, q1nc7) for complete human verification, comprising 504 reference expressions. We compute micro-averaged accuracy and scores for each binary attribute by comparing LLM predictions against the human gold standard. The human gold standard was collected by one experienced annotator (one of the paper authors) following the same annotation guidelines, with ambiguous cases discussed and resolved with other authors.
@@
The LLM annotations exhibit relatively high validity across multiple evaluation dimensions when compared to human gold standard. For all grounded reference expressions, the LLM achieves 95.5% accuracy and 99.5% micro-F1, demonstrating reliable recovery of each participant’s intended and interpreted referents. Table 2 then presents the micro-averaged evaluation results for the five binary attributes across all 3 dialogues. The LLM achieves high performance across all attributes, with accuracy ranging from around 97% to 100% and F1 scores from 0.89 to 1.00. At the reference expression level, 28 of 504 REs (5.6%) contained at least one attribute error, totaling 48 attribute-level mismatches. Attribute-level error breakdown shows that is_grounded has the highest error count (17), largely stemming from false negatives where the RE is grounded in gold by missed by the model, which typically occur when grounding is inferred from implicit confirmation or reuse rather than explicit echoing.
@@ 5.1. Classification of Understanding States
The addressee grounds the reference expression to a different landmark than the speaker intended. This is the critical case: both participants believe successful reference has occurred, yet their interpretations diverge. Such silent misalignments may persist undetected unless subsequent dialogue reveals the discrepancy.
@@
The interlocutors await further information or signal comprehension failure. This category includes several subtypes corresponding to communication breakdown scenarios associated with the attributes discussed in Section 4.1 (excluding is_imagined).
Table 3 presents the distribution of understanding states across our annotated corpus with the “misunderstood” type reaching 7.07%. However, as discussed in Section 3.2, 7.5% of landmarks exhibit lexical discrepancies where the same landmarks are labeled with different names on maps. For instance, when the giver says “the old mill” (referring to m12_old_mill@g) and the follower only has m12_mill_wheel@f on his/her map, they may initially correct each other but usually quickly reach an agreement that they are referring to the “same” landmark at the same location. In such cases, despite the initial lexical mismatch, we assume they have reached an aligned understanding state.
Verified by our observations, we unified the 10 pairs of lexical variants and convert the “misunderstood” cases caused by them into “Aligned” ones. Table 3 demonstrates the understanding state distribution after this unification as well. Now the misunderstanding rate drops to 1.82%, a significant reduction from the original 7.07%.
@@
Table 4 demonstrates how these misunderstandings distribute across different types of landmark discrepancies. The distribution shows a pattern: multiplicity discrepancies account for 50.9% of all misunderstandings (115/226 cases) despite representing only 7.3% of reference expressions in the corpus. The misunderstanding rate of this type is more then 6x higher than the corpus average (12.0% versus 1.86%). This reveals the success of the MapTask asymmetric design in eliciting misunderstandings, as multiplicity discrepancies inherently create ambiguity about which specific landmark instance is being referred to: when a landmark type appears twice on the giver’s map but only once on the follower’s map, participants often assume referential uniqueness Lane et al. (2006); Bard et al. (2000). For example, the giver refers to instance #0, the follower accommodates to their visible instance #1, and both believe successful reference has occurred. This silent divergence persists unless subsequent spatial descriptions reveal the mismatch.
@@
5.2. Understanding State Transition
To analyze how understanding evolves across repeated references to the same landmark, we extracted reference chains—sequences of reference expressions targeting the same landmark mtlm within a dialogue. Across 128 dialogues, we identified 1,665 reference chains with a mean length of 6.89 references per landmark (range: 1–70).
@@ 5.2. Understanding State Transition
| Multiplicity | 108 | 7.0 | 26 | 5 |
The length of these chains partly reflects how persistently participants negotiate understanding of landmarks and use these landmarks to generate spatial descriptions. Table 5 summarizes chain lengths by discrepancy type. Lexical variants yield the longest chains on average (10.7 references), as participants often need multiple turns to resolve naming mismatches and confirm shared understanding. Landmarks with no discrepancy generate the second longest chains, while existence discrepancies yield the shortest. Combined with our observations, this suggests that existence discrepancies are usually quickly ruled out as reference anchors when participants realize the landmark is absent from one map, whereas references to non-discrepant landmarks persist longer because they support iterative spatial description and navigation coordination, aligning with Viethen et al. (2011a); Reitter and Moore (2014)’s findings.
To further disentangle the efforts participants invest in coordinating understanding of landmarks from using grounded landmarks for navigation, we measure the number of turns (counted per utterance) between a chain’s first reference and its first aligned state (turns-to-ground). Figure 3 shows the cumulative distribution of turns-to-ground by discrepancy type across 1,334 chains that reached alignment. The longer negotiation for multiplicity likely reflects the cognitive complexity of distinguishing between multiple same-named landmarks versus simply confirming/denying existence. Participants must coordinate not only on the landmark name but also on which specific instance is intended, requiring more iterative spatial descriptions and confirmations.
There are 331 chains that hold no aligned state at all, indicating that participants sometimes stop early and abandon negotiation on some landmarks. This aligns with task-oriented dialogue efficiency.
@@ 5.3. Special Case
In the annotation, the LLM fails to capture the misalignment since the beginning, partly because of an omniscient/perspective bias: the model can “see” both maps at once, so it underestimates asymmetric knowledge and treats contradictions as noise rather than evidence of a wrong starting point. Prior Studies show that human addressees can achieve more efficient communication than overhearers, while VLMs perform poorly at overhearing Schober and Clark (1989); Wang et al. (2025). Future evaluations could explicitly test omniscience and perspective (analyzer versus participant versus overhearer conditions) to test models’ perspective-taking ability in grounded dialogue.
@@ 6. Conclusion
Our quantitative analyses reveal given the rarity of full misunderstandings, multiplicity discrepancies act as a consistent hazard for eliciting misalignment. Our statistics also support prior studies on interlocutors’ preferences for reusing grounding-established REs Healey et al. (2018); Viethen et al. (2011a); Reitter and Moore (2014).
We plan to establish evaluations based on these annotated dialogues for evaluating (V)LLM’ incremental grounding abilities under different information settings, moving beyond object identification toward perspective-taking in collaborative dialogue.
@@ 7. Limitations
While the LLM annotations achieves high performance compared to human gold standard (5.6% RE-level error rate; see Section 4.3), comprehensive inter-annotator agreement studies are needed to validate annotation reliability across diverse annotators.

### v2

@@ Abstract
| Nan Li1, Albert Gatt1, Massimo Poesio1,2 |
| 1Utrecht University, Utrecht, the Netherlands |
| 2Queen Mary University of London, London, the UK |
@@ 1. Introduction
RQ1 Is it possible to develop a personal-interpretation annotation scheme to capture nuanced understanding states for REs in collaborative tasks like MapTask?
@@
We operationalize reference expression grounding as a graded, perspective-dependent alignment process, and propose an annotation scheme at the RE level.
@@ 2. Related work
Formal dialogue theories have modeled interlocutors’ intentions and repairs through structured information state models. Some proposals Poesio and Traum (1997); Matheson et al. (2000) extended Discourse Representation Theory to integrate semantic and pragmatic information, showing how turn-taking, discourse segmentation, and grounding are represented through conversation acts in common ground. Ginzburg (2012) proposed the KoS framework111A term loosely connected to conversation-oriented semantics Ginzburg (2012)., where unresolved Questions Under Discussion (QUDs) persist until explicitly addressed. Lascarides and Asher (2009) modeled agreement and correction dynamics but assumed a highly idealized scenario which excludes misunderstandings. Schlangen and Skantze (2011) treated dialogue processing as an incremental hypothesis-formation and revision process operating at a sub-utterance level. Some recent work leveraged annotations from other meaning represenatations, e.g., Abstract Meaning Representation, to represent common ground structure and information states Khebour et al. (2024); Lai et al. (2025). We are informed by this line of work, but focus on RE-level perspectivist interpretations and inferring understanding states through the alignment and misalignment of these interpretations.
@@ 3.1. Development and Follow-up Studies
HCRC MapTask is a corpus documenting unscripted dialogues in a collaborative route-following task with asymmetric maps. Participants must collaborate to reproduce a route on one participant’s map (the instruction giver’s) on the other’s (the instruction follower’s). They cannot see each other’s maps and can only communicate verbally, with eye contact possible in half of the dialogues. The intentional differences between maps on routes, landmark names, and landmark placements often lead to ambiguous or misleading instructions. This makes the MapTask corpus a valuable resource and testbed for studying grounding, reference, and repair theories in dialogue.
Subsequent research analyzed reference expressions in MapTask from multiple perspectives. Bard et al. (2000) show that speakers reduce articulatory clarity when mentioning landmarks for the second time, even when the addressee has not heard the first mention or cannot see the referent, suggesting that intelligibility control is driven primarily by the speaker’s own knowledge rather than careful modeling of the listener’s information state. Varges (2005) treated spatial descriptions in MapTask as reference expressions that distinguish particular map points from distractors, observing that speakers frequently use spatial relations (e.g., “above the west lake”) and vague quantifiers (e.g., “about two inches southwest”) when referring to featureless route points around named landmarks. Marge and Rudnicky (2011) analyzed miscommunication patterns in MapTask, documenting how instruction followers detect ambiguity (e.g., multiple possible referents) and misunderstanding through clarification questions, repair strategies, and explicit feedback when requested actions appear impossible or unclear. Poesio et al. (2004) applied the MATE annotation scheme Poesio et al. (1999) to an Italian adaptation of MapTask to make anaphoric relations explicit.
@@ 4.1. Approach and General Principles
ordinal: Positional index (0, 1, …) for landmarks of the same type occurring multiple times on maps, ordered from bottom to top. Landmarks of this type have the same ordinal number if they are at the same location on both maps, which means it is possible that the only “parked van” on the follower’s map is m0_parked_van#1@f if the giver’s map has two “parked van”, while on the giver’s map m0_parked_van#0@g denotes the unique instance that is at a lower place.
@@ 4.2. Annotation Pipeline
The prompt (see Figure 2 and A.1) guides the model through a sequential and heuristic annotation workflow. First, the model identifies the speaker and addressee roles (giver/follower) and assigns the speaker’s intended landmark ID. Second, it evaluates is_quantificational to determine whether the expression presupposes a referent. If not quantificational, the model proceeds through the remaining four attributes in order. Finally, if is_grounded=true, the model assigns the addressee’s interpreted landmark ID. The model is asked to give a concise textual reason citing dialogue evidence for our further analysis.
@@ 4.3. Quality Evaluation
We apply the annotation pipeline to all 128 dialogues in the MapTask corpus, resulting in 13,077 annotated reference expression instances 555The LLM omitted annotations for 4 REs (from 3 different dialogues), despite their presence in both the dialogue context and the RE list, likely due to uncertainty during the generation process..
To evaluate the reliability of the annotation pipeline, we selected 3 dialogues (dialogue IDs: q1ec2, q1nc3, q1nc7) for complete human verification, comprising 504 reference expressions. We compute micro-averaged accuracy and f1-scores for each binary attribute by comparing LLM predictions against the human gold standard. The human gold standard was collected by one experienced annotator (one of the paper authors) following the same annotation guidelines, with ambiguous cases discussed and resolved with other authors.
@@
The LLM annotations exhibit relatively high validity across multiple evaluation dimensions when compared to human gold standard. For all grounded reference expressions, the LLM achieves 95.5% accuracy and 99.5% micro-F1, demonstrating reliable recovery of each participant’s intended and interpreted referents. Table 2 then presents the micro-averaged evaluation results for the five binary attributes across all 3 dialogues. The LLM achieves high performance across all attributes, with accuracy ranging from around 97% to 100% and F1 scores from 0.89 to 1.00. At the reference expression level, 28 of 504 REs (5.6%) contained at least one attribute error, totaling 48 attribute-level mismatches. Attribute-level error breakdown shows that is_grounded has the highest error count (17), largely stemming from false negatives where the RE is grounded in gold but missed by the model, which typically occur when grounding is inferred from implicit confirmation or reuse rather than explicit echoing.
@@ 5.1. Classification of Understanding States
The addressee grounds the reference expression to a different landmark from that the speaker intended. Both participants believe successful reference has occurred, yet their interpretations diverge. Such silent misalignments may persist undetected unless subsequent dialogue reveals the discrepancy.
@@
The interlocutors await further information or signal comprehension failure. This category includes several communication breakdown scenarios corresponding to the attributes discussed in Section 4.1, except for is_imagined.
For each RE, we infer the interlocutors’ understanding state from our annotated data, and Table 3 presents the distribution. The “misunderstood” type reaches 7.07%. However, as discussed in Section 3.2, 7.5% of landmarks exhibit lexical discrepancies where the same landmarks are labeled with different names on maps. For instance, when the giver says “the old mill” (referring to m12_old_mill@g) and the follower only has m12_mill_wheel@f on his/her map, they may initially correct each other but usually quickly reach an agreement that they are referring to the “same” landmark at the same location. In such cases, despite the initial lexical mismatch, we assume they have reached an aligned understanding state.
Verified by our observations, we unified the 10 pairs of lexical variants and adjust the “misunderstood” cases caused by them into “Aligned” ones. Table 3 demonstrates the understanding state distribution after this unification as well. Now the misunderstanding rate drops to 1.82%, a significant reduction from the original 7.07%.
@@
Table 4 demonstrates how these misunderstandings distribute across different types of landmark discrepancies. The distribution shows a pattern: multiplicity discrepancies account for 48.5% of all misunderstandings (115/239 cases) despite representing only 7.3% of reference expressions in the corpus. The misunderstanding rate of this type is more than 6 higher than the corpus average (12.0% versus 1.8%). This reveals the success of the MapTask asymmetric design in eliciting misunderstandings, as multiplicity discrepancies inherently create ambiguity about which specific landmark instance is being referred to: when a landmark type appears twice on the giver’s map but only once on the follower’s map, participants often assume referential uniqueness Lane et al. (2006); Bard et al. (2000). For example, the giver refers to instance #0, the follower accommodates to their visible instance #1, and both believe successful reference has occurred. This silent divergence persists unless subsequent spatial descriptions reveal the mismatch.
@@
5.2. Reference Chain
To analyze how understanding evolves across repeated references to the same landmark, we extracted reference chains—sequences of reference expressions targeting the same landmark (same mtlm) within a dialogue. Across 128 dialogues, we identified 1,665 reference chains with a mean length of 6.89 REs per landmark (range: 1–70).
@@ 5.2. Reference Chain
| Multiplicity | 108 | 7.00 | 26 | 5 |
The length of these chains partly reflects how persistently participants negotiate understanding of landmarks and use these landmarks to generate spatial descriptions. Table 5 summarizes chain lengths by discrepancy type. Lexical variants yield the longest chains on average (10.7 references), as participants often need multiple turns to resolve naming mismatches and confirm shared understanding. Landmarks with no discrepancy generate the second longest chains on average, while existence discrepancies yield the shortest. This suggests that existence discrepancies are usually quickly ruled out as reference anchors when participants realize the landmark is absent from one map, whereas references to non-discrepant landmarks persist longer because they support iterative spatial description and navigation coordination, aligning with previous findings Viethen et al. (2011a); Reitter and Moore (2014).
To illustrate how understanding states evolve through a reference chain, we present a simple multiplicity discrepancy case from dialogue q7ec2. The landmark vast meadow appears twice on the giver’s map (m11_vast_meadow#0@g, lower/unique; m11_vast_meadow#1@g, upper/shared) but only once on the follower’s (m11_vast_meadow#1@f, upper/shared).
Table 6 traces five reference expressions to this landmark (see Figure A.2 in Appendix A.2 for the dialogue excerpt). The giver initiates with “a meadow” (ref.112), intending the lower instance (#0@g). The follower queries “a vast meadow?” (ref.114) and the giver confirms (ref.115). However, the follower identifies his/her own, the upper shared, vast meadow (#1@f), creating a misunderstanding: both have landmarks in mind, yet they are different ones.
| Ref | Spk | Expression | Giver ID | Follower ID | State |
| 112 | G | “a meadow” | #0@g | ungrounded | Pending |
| 114 | F | “a meadow” | #0@g | ungrounded | Pending |
| 115 | G | “a vast meadow” | #0@g | #1@f | Misunderstood |
| 116 | F | “a vast meadow” | #1@g | #1@f | Aligned |
| 119 | F | “vast meadow” | #1@g | #1@f | Aligned |
| 127 | G | “a vast meadow” | #0@g | ungrounded | Pending |
In the next turn (ref.116), the follower describes the location as “just below the blacksmiths and carpenters,” which corresponds to the upper meadow shared on both maps. The giver accepts this reframing via confirmation, and both participants align on the upper instance (#1). This alignment persists through ref.119. However, later in the dialogue (ref.127), the giver again mentions “a vast meadow” in a different spatial context, referring back to the lower instance (#0@g), but the follower cannot ground it, reverting the state to pending.
This chain exemplifies how multiplicity discrepancies create recurring alignment challenges: the same landmark name maps to different physical locations on the maps for participants, producing misunderstandings that may resolve through spatial negotiation but can resurface later in the dialogue.
5.3. Turns to Ground
To further disentangle the efforts participants invest in coordinating understanding of landmarks from using grounded landmarks for navigation, we measure the number of turns (counted per utterance) between a chain’s first reference and its first aligned state, and term this measure turns-to-ground. Figure 3 shows the cumulative distribution of turns-to-ground by discrepancy type across 1,334 chains that reached alignment. The longer negotiation for multiplicity likely reflects the cognitive complexity of distinguishing between multiple same-named landmarks versus simply confirming/denying existence. Participants must coordinate not only on the landmark name but also on which specific instance is intended, requiring more iterative spatial descriptions and confirmations.
There are 331 chains that never reach an aligned state at all, indicating that participants sometimes stop early and abandon negotiation on some landmarks. This aligns with task-oriented dialogue efficiency.
@@ 5.4. Special Case
In the annotation, the LLM fails to capture the misalignment since the beginning, partly because of an omniscient bias: the model can “see” both maps at once, so it underestimates asymmetric knowledge and treats contradictions as noise rather than evidence of a wrong starting point. Prior studies show that human addressees can achieve more efficient communication than overhearers, while VLMs perform poorly at overhearing Schober and Clark (1989); Wang et al. (2025). Future evaluations could explicitly test omniscience and perspective (analyzer versus participant versus overhearer conditions) to test models’ perspective-taking ability in grounded dialogue.
@@ 6. Conclusion
Our quantitative analyses reveal that given the rarity of full misunderstandings, multiplicity discrepancies act as a consistent hazard for eliciting misalignment. Our statistics also support prior studies on interlocutors’ preferences for reusing grounding-established REs Healey et al. (2018); Viethen et al. (2011a); Reitter and Moore (2014).
We plan to establish evaluations based on these annotated dialogues for evaluating (V)LLMs’ incremental grounding abilities under different information settings, moving beyond object identification toward perspective-taking in collaborative dialogue.
@@ 7. Limitations
Our annotation pipeline relies on a single LLM (GPT-5), and the human gold standard was produced by one annotator (with ambiguous cases discussed among co-authors). This constrains claims about cross-model generalizability and inter-annotator agreement. Although our annotation scheme is desingned as a structured and heuristic decision cascade with JSON-shema validation and explicit explanations — achieving 94.4% RE-level accuracy across 504 reference expressions (see Section 4.3) — these results reflect one model under one set of conditions. Comprehensive evalution across multiple LLMs and human annotators remains an important direction for future work.
@@
8. Ethics Statement
This research involves annotated dialogue data from the HCRC MapTask corpus, which consists of recordings of human participants engaged in a collaborative task. The corpus is publicly available for research purposes and doesn’t contain any sensitive personal information.
9. Code and Data Availability
We release our full code, annotation prompt templates, and the complete annotated dataset to facilitate reproducibility666https://github.com/chnln/grounded-misunderstandings-in-maptask.
10. Acknowledgements
We appreciate the helpful comments and suggestions from the anonymous reviewers. We would also like to thank Amy Isard, Matthew Aylett, and Martín Villalba for their valuable input on the analysis of misunderstanding in MapTask during the early stages of this project. This work is funded by the Dutch Research Council (NWO) through the AiNed Fellowship Grant NGF.1607.22.002, Dealing with Meaning Variation in NLP.
Appendix A Appendices
A.1. Annotation Prompt Template
Figure 2 in the main text shows an abbreviated version of the annotation prompt template. Below we provide the full content of each section in the prompt template777To enforce the model’s output to adhere to the required format, we not only include the output format section in the prompt, but also attach a JSON-schema file as output format parameter when calling the OpenAI Response API.. Each section is wrapped in the XML style in the actual prompt, for example, the background section is enclosed by <background> and </background>.
Figure 4: Full annotation prompt template.
A.2. Reference Chain Dialogue Excerpt
Table 6 in the main text traces the understanding state transitions for the landmark vast meadow in dialogue q7ec2. Below is the relevant dialogue excerpt providing context for those reference expressions (refs 112, 114, 115, 116, 119, 127). Reference expressions are enclosed in <<...>>; the six reference expressions are bolded.
[follower utt:132] <<a vast meadow>>ref.114
[giver utt:133] <<a vast meadow>>ref.115
[follower utt:134] i’ve got <<a vast meadow>>ref.116 up just below <<the blacksmiths>> and <<carpenters>>
[giver utt:135] well is that huge big and there’s nothing at all in that space
[follower utt:136] uh-huh except
[giver utt:137] about halfway up the page
[follower utt:138] <<vast meadow>>ref.119 about halfway up the page it’s empty on mine
[giver utt:139] well you want to well anyway follow <<crane bay>> round <<the curve>>
… (utterances 140–160: participants navigate along crane bay and attractive cliffs) ...
[giver utt:161] just below i’ve got <<a vast meadow>>ref.127 there
[follower utt:162] right where will i stop that line across the top
Figure 5: Dialogue excerpt from q7ec2 (utterances 131–161).
Context: The giver’s map contains two vast meadow instances: #0@g (lower, unique to giver) and #1@g (upper, shared with follower). The follower has only the upper instance (#1@f). At ref.112, the giver refers to the lower meadow. The follower’s “a vast meadow?” (utt:132) is a clarification query. When the giver confirms (ref.115), the follower identifies the upper shared meadow instead, producing a misunderstanding. At ref.116, the follower’s spatial description (“below the blacksmiths and carpenters”) shifts the giver’s grounding target to the upper/shared meadow, and both align. After a 28-utterance gap discussing other landmarks, the giver returns to “a vast meadow” (ref.127) in a new spatial context, but the follower is not able to ground it — the state reverts to pending.